# Direct Video-Based Spatiotemporal Deep Learning for Cattle Lameness Detection

### v3

@@ Abstract
keyword1, Keyword2, Keyword3, Keyword4
@@ 1 Introduction
Livestock production is a cornerstone of global food security, and the dairy and beef sectors depend critically on healthy, mobile animals. Lameness—typically arising from hoof lesions, interdigital dermatitis, or gait-related injuries—remains one of the most prevalent welfare and economic challenges in cattle farming. A recent systematic review by [1] estimates that 23–25% of dairy cows worldwide exhibit clinical or sub-clinical lameness at any given time, leading to reduced milk yield, delayed reproduction, and increased culling rates. Because lameness pain alters key behaviours such as walking, standing, and feeding, early detection is essential for prompt treatment and cost containment.
@@
    End-to-end deep learning architecture. We design and compare two spatiotemporal models—3D Convolutional Neural Networks (3D-CNN) and Convolutional Long Short-Term Memory networks (ConvLSTM2D)—that operate directly on video tensors without intermediate object-detection or pose-estimation steps.
@@ 2 Related Work
Cattle farming plays a crucial role in ensuring global food security by supplying essential meat and dairy products. With rising global demand, large-scale cattle farming has become increasingly common. However, this trend introduces operational challenges, particularly in monitoring animal health and welfare, which are directly tied to productivity outcomes [2, 3, 25]. One of the most prevalent and costly health issues in cattle is lameness—a painful condition often linked to gait-related injuries or hoof disorders. Studies estimate that approximately one in four cattle are affected by lameness at any given time, resulting in reduced productivity and higher production costs, which can contribute to elevated market prices for meat and dairy [4, 6, 5].
@@
Both manual and automated CV-based approaches rely on observing behavioral cues to identify lameness. Lameness typically manifests as altered locomotion, reduced activity, or abnormal postures—behaviors that can be detected during routine activities like walking, standing, or feeding. Modern deep learning models can analyze video data from multiple perspectives to capture such behavioral cues and distinguish between normal and abnormal gait patterns [16, 17]. These models focus on learning the spatial and temporal dynamics of movement, allowing automated systems to detect early deviations and classify them accordingly.
@@
Additional approaches have explored object detection and pose estimation pipelines. Barney et al. [20] proposed a modified Mask R-CNN architecture for real-time lameness detection using multi-cattle pose estimation. In a similar context, researchers in [19] combined object detection models (Mask R-CNN, YOLO) with post-hoc machine learning classifiers. Each detected cow was assigned a local identifier for tracking, followed by feature extraction and classification. Russello et al. [6] designed a deep learning model that extracted nine locomotion traits from videos and used them to detect lameness. In their study, Wu et al. [19] used 50 cattle videos—20 lame and 30 non-lame—and randomly sampled 5,000 frames, focusing on leg movement. Their LSTM model achieved a classification accuracy of 98.57%, outperforming traditional ML baselines under a 50:50 train–test split.
Despite promising results, many of these studies relied on data captured in restricted, controlled farm environments, often using small sample sizes and homogeneous conditions. The generalizability of such models to real-world settings, where lighting, surface conditions, and animal orientation vary, remains limited. Furthermore, most existing pipelines involve multiple processing stages—such as object detection, tracking, and pose estimation—which increase complexity and computational overhead.
@@
We then trained two spatiotemporal deep learning models—3D CNN and ConvLSTM2D—on this dataset and evaluated their performance using standard classification metrics. The results demonstrate that our approach, which avoids multi-step feature engineering, can effectively classify cattle lameness with performance comparable to or exceeding that of existing state-of-the-art models. Our findings highlight the potential of internet-sourced video data and simplified model pipelines for scalable livestock health monitoring in real-world scenarios.

### v4

@@ Abstract
Cattle, Lameness detection, Deep-learning, Computer vision techniques, Image processing
@@ 1 Introduction
Livestock production is a cornerstone of global food security, and the dairy and beef sectors depend critically on healthy, mobile animals. Lameness, typically arising from hoof lesions, interdigital dermatitis, or gait-related injuries remains one of the most prevalent welfare and economic challenges in cattle farming. A recent systematic review by [1] estimates that 23–25% of dairy cows worldwide exhibit clinical or sub-clinical lameness at any given time, leading to reduced milk yield, delayed reproduction, and increased culling rates. Because lameness pain alters key behaviours such as walking, standing, and feeding, early detection is essential for prompt treatment and cost containment.
@@
    End-to-end deep learning architecture. We design and compare two spatiotemporal models—3D Convolutional Neural Networks (3D-CNN) and Convolutional Long Short-Term Memory networks (ConvLSTM2D), that operate directly on video tensors without intermediate object-detection or pose-estimation steps.
@@ 2 Related Work
Cattle farming plays a crucial role in ensuring global food security by supplying essential meat and dairy products. With rising global demand, large-scale cattle farming has become increasingly common. However, this trend introduces operational challenges, particularly in monitoring animal health and welfare, which are directly tied to productivity outcomes [2, 3, 25]. One of the most prevalent and costly health issues in cattle is lameness, a painful condition often linked to gait-related injuries or hoof disorders. Studies estimate that approximately one in four cattle are affected by lameness at any given time, resulting in reduced productivity and higher production costs, which can contribute to elevated market prices for meat and dairy [4, 6, 5].
@@
Both manual and automated CV-based approaches rely on observing behavioral cues to identify lameness. Lameness typically manifests as altered locomotion, reduced activity, or abnormal postures behaviors that can be detected during routine activities like walking, standing, or feeding. Modern deep learning models can analyze video data from multiple perspectives to capture such behavioral cues and distinguish between normal and abnormal gait patterns [16, 17]. These models focus on learning the spatial and temporal dynamics of movement, allowing automated systems to detect early deviations and classify them accordingly.
@@
Additional approaches have explored object detection and pose estimation pipelines. Barney et al. [20] proposed a modified Mask R-CNN architecture for real-time lameness detection using multi-cattle pose estimation. In a similar context, researchers in [19] combined object detection models (Mask R-CNN, YOLO) with post-hoc machine learning classifiers. Each detected cow was assigned a local identifier for tracking, followed by feature extraction and classification. Russello et al. [6] designed a deep learning model that extracted nine locomotion traits from videos and used them to detect lameness. In their study, Wu et al. [19] used 50 cattle videos (20 lame and 30 non-lame) and randomly sampled 5,000 frames, focusing on leg movement. Their LSTM model achieved a classification accuracy of 98.57%, outperforming traditional ML baselines under a 50:50 train–test split.
Despite promising results, many of these studies relied on data captured in restricted, controlled farm environments, often using small sample sizes and homogeneous conditions. The generalizability of such models to real-world settings, where lighting, surface conditions, and animal orientation vary, remains limited. Furthermore, most existing pipelines involve multiple processing stages—such as object detection, tracking, and pose estimation which increase complexity and computational overhead.
@@
We then trained two spatiotemporal deep learning models, 3D CNN and ConvLSTM2D on this dataset and evaluated their performance using standard classification metrics. The results demonstrate that our approach, which avoids multi-step feature engineering, can effectively classify cattle lameness with performance comparable to or exceeding that of existing state-of-the-art models. Our findings highlight the potential of internet-sourced video data and simplified model pipelines for scalable livestock health monitoring in real-world scenarios.